### v1

@@ Abstract
Several defense mechanisms have been proposed to sanitize training data-sets. However, these methods often rely on the availability of a clean dataset to compute statistical anomalies, which may not always be feasible in real-world scenarios where datasets can be unavailable or compromised. To address this limitation, we propose a novel defense method that trains a model on the given dataset, detects poisoned classes, and extracts the critical part of the attack trigger before identifying the poisoned instances. This approach enhances explainability by explicitly revealing the harmful part of the trigger. The effectiveness of our method is demonstrated through experimental evaluations on well-known image datasets and comparative analysis against three state-of-the-art algorithms: SCAn, ABL, and AGPD.
@@ 1 Introduction
The rise of Artificial Intelligence (AI), and more specifically Deep Learning-based systems, has been dazzling and unparalleled. Its use is now widespread and, thanks to the availability of free software, a large public, even without specific knowledge, can build an AI model. They can also download training datasets and pre-trained models. Like any software, AI systems are vulnerable to malicious attacks. They can traditionally be attacked at the software layer, e.g., through vulnerabilities in the software libraries, but they can also be attacked through data and model-parameter manipulation. A corrupted AI can exhibit malicious behavior, which can have critical consequences. For example, self-driving cars may intentionally collide depending on the attacker’s aim and a Large Language Model (LLM) may give fake news. This security risk has become a real-life threat and a major issue when researchers have shown that the AI models can be attacked in various way depending on the aim, knowledge, and model access privilege of the attacker [16, 21]. The attacker can lead an evasion attack by slightly modifying the model’s input in order to change the decision/prediction. For example, for a spam recognition system, the attacker may change some words to bypass the filter. She can also lead inference attacks where the attacker wants to extract some information about the training dataset on which the model was trained. This attack is also known as privacy attack and model inversion attack. Indeed, in a personalized drug system, she may want to know if a particular patient information was in the training set. During the training stage, an attacker can also poison the training data to either degrade the model’s performance (non-targeted attacks), e.g., by flipping the labels, or introduce a backdoor than can be exploited, during inference, to trigger a malicious behavior of the model.
@@
In this paper, we focus on a defense method against the backdoor attacks where the attacker wants a model to associate her data pattern to her target label. Thus, in the inference stage, any input containing the pattern will trigger the misclassification of the input as the target label. This is a vicious attack as the model seems to be pristine w.r.t. to its performance on clean data but in fact contains a backdoor which changes the prediction only in the presence of the trigger in the input data. For example, in a self-driving car, an attacker may have corrupted the model to recognize a green-square sticker pasted on a stop board as a 70mph speed limit board. Such attacks can easily be performed when the attacker has a write-access to the training data, as illustrated by the BadNets algorithm [7]. More advanced attacks based on BadNets have been proposed, but they often require more resources or access rights. By its simplicity, BadNets remains a threat that can be exploited by a large population, including opportunistic criminals and script-kiddies (open source script users without particular knowledge). Many defenses have been proposed in the literature, but most of them require an additional dataset of clean data on which statistics are computed to detect a distribution anomaly induced by a poisoned sample. In real-life, such dataset may hardly be available and even be subject to a corruption.
@@
Our contributions can be summarized in three points: 1) we propose a new defense algorithm against BadNets [7] and Blended [2] attacks which does not rely on additional prerequisites such as the availability of a clean dataset, 2) the defense is explainable by the fact that we extract the important part of the trigger that causes the malicious behavior, and 3) our method works in All-to-All attacks where many classes are poisoned.
@@ 3.3 Method
The proposed data sanitization method, described in algorithm 1, is composed of 7 main steps:
@@
In the following paragraphs, we describe each steps.
@@ 3.3.2 Stage 2: Poisoned classes detection
For poisoned instances, some components of the gradient (Eq. 5) will have a large absolute value compared to the clean-instances. Indeed, the partial derivative at given coordinates tells us how much the loss function will change if the pixel changes by one unit. We can then assume that, for poison instances, a few pixels will have a large impact on the model’s output. Thus, the most important pixel/variable is located at the indexes $(i^{*},j^{*})$ (after taking the mean of the gradient over the color channel):
@@ 3.3.4 Stage 4: Check and filter the patterns
At this stage, a candidate pattern and its mask have been calculated for each potentially poisoned class. We need to check that it triggers a malicious behavior of the simple poisoned model. For that, we patch the pattern to instances from other classes (estimated ${\tilde{\mathcal{D}_{C}}}$) before feeding them to the model, and we check if the model assign them to the class associated with the pattern. We compute the Attack Success Rate (ASR), i.e., the proportion of patched instances predicted as the attack label. If this rate falls below a user-defined threshold, we remove the pattern from the candidate set. The remaining patterns are then considered harmful.
@@
Figure 2: Illustration of the candidate pattern computation process. The first row shows (2(a)) an example of input image from the cat class, (2(b)) its malicious version labeled with the attacker’s label and (2(c)) its gradient (eq. 5). In the last two rows, the pattern computation process, described in the algorithm 2, is illustrated step by step. The gray background in (2(j)) represents the transparent color.
@@ 4.1 Experimental Setup
For the evaluation, we used the BackdoorBench framework [26], and our code is available online [^1] https://github.com/Orange-OpenSource/BackdoorBench/tree/vbd-v1. We compared our method, VBD, against three high-performing state-of-the-art methods [28]: Anti-Backdoor Learning (ABL) [14], AGPD [28], and SCAn [25]. For VBD, we used an ensemble of 3 models: a 1-layer model with 64 filters, a 2-layers model with 10 filters each, and a 2-layers model with 16 filters each, following the architecture shown in 1. We used 80% of the training set only to train the models and the remaining 20% for the validation, detection, pattern extraction, and detector training. For the competitors, we used the default settings, which include the use of the Pre-Act ResNet18 model [9], knowing that AGPD and SCAn require also an additional dataset of clean images, which, by default, consists of 10 images per class taken from the test set.
@@
We used six attack triggers to evaluate the methods, as shown in Figure 3. When more than one class is poisoned (All-to-All), we used only the grid and square triggers, positioning them at different locations based on the class index. The placement starts at the bottom right of the image and moves from right to left and bottom to top to avoid overlapping. We ran a total of 80 experiments: for both CIFAR-10 and Tiny ImageNet, there were 24 experiments (6 triggers × 4 blending factors) for the All-to-One setting and 16 experiments (4 triggers $\times$ 4 blending factors) for the All-to-All case. Each experiment is repeated 5 times. In each repetition $i\in\{0,\ldots,4\}$, the random seed is set to $i$, and for the All-to-One attacks, the target class is also set to $i$. The F1-scores and the averaged F1-scores are reported as the mean over the five repetitions, along with their standard deviation.
@@ 4.2 Experimental Results
Table 1 provides the F1-score of the defense methods for the All-to-One BadNets attack. SCAn and VBD achieve the best results. On average, SCAn outperforms VBD by $0.34\%$ on CIFAR-10, while VBD performs better on Tiny ImageNet, with an improvement of $1.59\%$. The average F1-score across the 12 experiments indicates that VBD surpasses SCAn by $0.62\%$. Notably, AGPD exhibits low detection performance on the "Big Kitty" pattern, as this pattern covers the entire image, making it difficult for AGPD to distinguish between poisoned and clean instances.
Table 1: All-to-One BadNets poisoned-instance detection F1-score (%).
@@
For the All-to-All BadNets attacks, SCAn fails on both CIFAR-10 and Tiny ImageNet as shown in Table 2. It appears that when at least half of the classes are poisoned SCAn is unable to detect the attack. On the contrary, VBD achieves the best F1-performance, outperforming AGPD by $8.69\%$ on CIFAR-10 and ABL by $23\%$ on Tiny-ImageNet.
Table 2: All-to-All BadNets poisoned-instance detection F1-score (%).
@@
The performance results for the Blended attacks in both All-to-One and All-to-All settings are provided in Table 3 and Table 4, respectively. As we can see, in the All-to-One setting, the results are comparable to those in the All-to-One BadNets case, with SCAn and VBD leading. Nevertherless, in this case, VBD outperforms SCAn by $0.24\%$ on CIFAR-10 and by $7.37\%$ on Tiny ImageNet. The averaged F1 score across the experiments on both datasets shows that VBD outperforms SCAn by $3.8\%$. For the All-to-All Blended attack case, as previously mentioned, SCAn fails, and AGPD outperforms the other defenses, with VBD finishing second. It is noteworthy that VBD struggles to detect the attack and extract the triggers when the blending factor is below $50\%$. However, when the blending factor is $50\%$, VBD performs better than AGPD.
Table 3: All-to-One Blended poisoned-instance detection mean F1-score (%) over the 6 triggers shown in Figure 3.
@@
Table 4: All-to-All Blended poisoned-instance detection mean F1-score (%) over the 4 grid and square triggers shown in Figure 3.
@@
Table 5 provides a summary of the performance of the tested defense methods, with the F1-score averaged over all the 80 experiments we conducted. VBD ranks first on CIFAR-10 and, notably, on Tiny ImageNet. It outperforms AGPD by $3.61\%$ on CIFAR-10, and ABL by $6.25\%$ on Tiny ImageNet. SCAn is penalized by its failure in the All-to-All settings. It is worth noting that, unlike AGPD and SCAn, VBD and ABL do not rely on an auxiliary dataset. The standard deviation also shows that VBD is a stable method.
Table 5: Summary of the poisoned-instance detection F1-score (%) averaged over all the 80 experiments.

### v2

@@ Abstract
Several defense mechanisms have been proposed to sanitize training datasets. However, these methods often rely on the availability of a clean dataset to detect statistical anomalies, which may not always be feasible in real-world scenarios where datasets can be unavailable or compromised. To address this limitation, we propose a novel defense method that trains a model on the given dataset, detects poisoned classes, and extracts the critical part of the attack trigger before identifying the poisoned instances. This approach enhances explainability by explicitly revealing the harmful part of the trigger. The effectiveness of our method is demonstrated through experimental evaluations on well-known image datasets and a comparative analysis against three state-of-the-art algorithms: SCAn, ABL, and AGPD.
@@ 1 Introduction
The rise of Artificial Intelligence (AI), and more specifically Deep Learning-based systems, has been dazzling and unparalleled. Its use is now widespread and, thanks to the availability of free software, a large public, even without specific knowledge, can build an AI model. They can also download training datasets and pre-trained models. Like any software, AI systems are vulnerable to malicious attacks. They can traditionally be attacked at the software layer, e.g., through vulnerabilities in the software libraries, but they can also be attacked through data and model-parameter manipulation. A corrupted AI can exhibit malicious behavior, which can have critical consequences. For example, self-driving cars may intentionally collide depending on the attacker’s aim and a Large Language Model (LLM) may give fake news. This security risk has become a real-life threat and a major issue when researchers have shown that the AI models can be attacked in various ways depending on the aim, knowledge, and model access privilege of the attacker [16, 21]. The attacker can lead an evasion attack by slightly modifying the model’s input in order to change the decision/prediction. For example, for a spam recognition system, the attacker may change some words to bypass the filter. She can also lead inference attacks where the attacker wants to extract some information about the training dataset on which the model was trained. This attack is also known as privacy attack or model inversion attack. Indeed, in a personalized drug system, she may want to know if a particular patient information was in the training set. During the training stage, an attacker can also poison the training data to either degrade the model’s performance (non-targeted attacks), e.g., by flipping the labels, or introduce a backdoor than can be exploited, during inference, to trigger a malicious behavior of the model.
@@
In this paper, we focus on a defense method against the backdoor attacks where the attacker wants a model to associate her data pattern to her target label. Thus, in the inference stage, any input containing the pattern will trigger the misclassification of the input as the target label. This is a vicious attack as the model seems to be pristine w.r.t. to its performance on clean data but in fact contains a backdoor which changes the prediction only in the presence of the trigger in the input data. For example, in a self-driving car, an attacker may have corrupted the model to recognize a green-square sticker pasted on a stop sign as a 70mph speed limit board. Such attacks can easily be performed when the attacker has a write access to the training data, as illustrated by the BadNets algorithm [7]. More advanced attacks based on BadNets have been proposed, but they often require more resources or access rights. By its simplicity, BadNets remains a threat that can be exploited by a large population, including opportunistic criminals and script-kiddies (open source script users without particular knowledge). Many defenses have been proposed in the literature, but most of them require an additional dataset of clean data on which statistics are computed to detect a distribution anomaly induced by a poisoned sample. In real-life, such dataset may hardly be available and even be subject to a corruption.
@@
We propose a multi-stage defense algorithm that estimates the poisoned subset of the training data, detects the poisoned classes, extracts the attack trigger for each class, and identifies the poisoned instances based on the extracted pattern. Our contributions can be summarized in three points: 1) we introduce a novel defense algorithm against BadNets [7] and Blended [2] attacks which does not rely on additional prerequisites such as the availability of a clean dataset; 2) the defense is explainable, as it extracts the most important part of the trigger responsible for the malicious behavior; and 3) our method is effective in All-to-All attacks where many classes are poisoned.
@@ 3.3 Method
Figure 1: Workflow of the proposed sanitization method.
The proposed data sanitization method is described in Algorithm 1 and illustrated step by step in Figure 1. It consists of seven main steps, listed as follows:
@@
In the following paragraphs, we provide a detailed description of each step.
@@ 3.3.2 Stage 2: Poisoned classes detection
For poisoned instances, some components of the gradient (Eq. 5) will have a large absolute value compared to the clean-instances. Indeed, the partial derivative at given coordinates indicates the sensitivity of the loss function to changes in that pixel’s value. In the case of poisoned instances, the backdoor pattern is designed to strongly influence the model’s prediction. As a result, only a few pixels, those containing the trigger, tend to dominate the gradient signal. We thus assume that, for poisoned instances, a small number of pixels will have a large impact on the prediction. The most influential pixel can therefore be identified by locating the index $(i^{*},j^{*})$ corresponding to the maximum absolute value of the mean gradient across the color channels:
@@ 3.3.4 Stage 4: Check and filter the patterns
At this stage, a candidate pattern and its mask have been calculated for each potentially poisoned class. We need to check that it triggers a malicious behavior of the simple poisoned model. To this end, we patch the pattern to instances from other classes (estimated ${\tilde{\mathcal{D}_{C}}}$) before feeding them to the model, and we check whether the model assigns them to the class associated with the pattern. We compute the Attack Success Rate (ASR), i.e., the proportion of patched instances predicted as the attack label. If this rate falls below a user-defined threshold, we remove the pattern from the candidate set. The remaining patterns are then considered harmful.
@@
Figure 3: Illustration of the candidate pattern computation process. The first row shows (3(a)) an example of input image from the cat class, (3(b)) its malicious version labeled with the attacker’s label and (3(c)) its gradient (Eq. 5). In the last two rows, the pattern computation process, described in the Algorithm 2, is illustrated step by step. The gray background in (3(j)) represents transparency.
@@ 4.1 Experimental Setup
For the evaluation, we used the BackdoorBench framework [26], and our code is available online [^1] https://github.com/Orange-OpenSource/BackdoorBench/tree/vbd-v1. We compared our method, VBD, against three high-performing state-of-the-art methods [28]: Anti-Backdoor Learning (ABL) [14], AGPD [28], and SCAn [25]. For VBD, we used an ensemble of 3 models: a 1-layer model with 64 filters, a 2-layers model with 10 filters each, and a 2-layers model with 16 filters each, following the architecture shown in Figure 2. We used 80% of the training set only to train the models and the remaining 20% for the validation, detection, pattern extraction, and detector training. For training the poisoned and detector models, we used a learning rate of 0.01, a batch size of 256 for CIFAR-10 and 64 for Tiny ImageNet, and 20 training epochs. For the competitors, we used the default settings, which include the use of the Pre-Act ResNet18 model [9], knowing that AGPD and SCAn require also an additional dataset of clean images, which, by default, consists of 10 images per class taken from the test set. Full implementation details and additional parameters are available in the released code repository.
@@
We used six attack triggers to evaluate the methods, as shown in Figure 4. When more than one class is poisoned (All-to-All), we used only the grid and square triggers, positioning them at different locations based on the class index. The placement starts at the bottom right of the image and moves from right to left and bottom to top to avoid overlapping. We ran a total of 80 experiments: for both CIFAR-10 and Tiny ImageNet, there were 24 experiments (6 triggers × 4 blending factors) for the All-to-One setting and 16 experiments (4 triggers $\times$ 4 blending factors) for the All-to-All case. Each experiment is repeated 5 times. In each repetition $i\in\{0,\ldots,4\}$, the random seed is set to $i$, and for the All-to-One attacks, the target class is also set to $i$. The $F_{1}$-scores and the averaged $F_{1}$-scores are reported as the mean over the five repetitions, along with their standard deviation.
@@
4.2 Evaluation Metrics
To evaluate the performance of poisoned instance detection, we use the $F_{1}$ score, defined as follows:
$F_{1}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}},\hskip 11.38092pt\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}},\hskip 11.38092pt\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$
where TP, TN, FP, and FN denote the number of true positives, true negatives, false positives, and false negatives, respectively.
Precision measures the proportion of correctly identified poisoned instances among all instances classified as poisoned. Recall measures the proportion of actual poisoned instances that were correctly identified by the classifier. Since it is important for a method to perform well on both metrics, the $F_{1}$ score, defined as the harmonic mean of precision and recall, is used to provide a balanced evaluation.
@@ 4.3 Experimental Results
Table 1 provides the $F_{1}$-score of the defense methods for the All-to-One BadNets attack. SCAn and VBD achieve the best results. On average, SCAn outperforms VBD by $0.34\%$ on CIFAR-10, while VBD performs better on Tiny ImageNet, with an improvement of $1.59\%$. The average $F_{1}$-score across the 12 experiments indicates that VBD surpasses SCAn by $0.62\%$. Notably, AGPD exhibits low detection performance on the "Big Kitty" pattern, as this pattern covers the entire image, making it difficult for AGPD to distinguish between poisoned and clean instances.
Table 1: All-to-One BadNets poisoned-instance detection $F_{1}$-score (%).
@@
For the All-to-All BadNets attacks, SCAn fails on both CIFAR-10 and Tiny ImageNet as shown in Table 2. It appears that when at least half of the classes are poisoned SCAn is unable to detect the attack. On the contrary, VBD achieves the best $F_{1}$-performance, outperforming AGPD by $8.69\%$ on CIFAR-10 and ABL by $23\%$ on Tiny-ImageNet.
Table 2: All-to-All BadNets poisoned-instance detection $F_{1}$-score (%).
@@
The performance results for the Blended attacks in both All-to-One and All-to-All settings are provided in Table 3 and Table 4, respectively. As we can see, in the All-to-One setting, the results are comparable to those in the All-to-One BadNets case, with SCAn and VBD leading. Nevertheless, in this case, VBD outperforms SCAn by $0.24\%$ on CIFAR-10 and by $7.37\%$ on Tiny ImageNet. The averaged $F_{1}$ score across the experiments on both datasets shows that VBD outperforms SCAn by $3.8\%$. For the All-to-All Blended attack case, as previously mentioned, SCAn fails, and AGPD outperforms the other defenses, with VBD finishing second. It is noteworthy that VBD struggles to detect the attack and extract the triggers when the blending factor is below $50\%$. However, when the blending factor is $50\%$, VBD performs better than AGPD.
Table 3: All-to-One Blended poisoned-instance detection mean $F_{1}$-score (%) over the 6 triggers shown in Figure 4.
@@
Table 4: All-to-All Blended poisoned-instance detection mean $F_{1}$-score (%) over the 4 grid and square triggers shown in Figure 4.
@@
Table 5 provides a summary of the performance of the tested defense methods, with the $F_{1}$-score averaged over all the 80 experiments we conducted. VBD ranks first on CIFAR-10 and, notably, on Tiny ImageNet. It outperforms AGPD by $3.61\%$ on CIFAR-10, and ABL by $6.25\%$ on Tiny ImageNet. SCAn is penalized by its failure in the All-to-All settings. It is worth noting that, unlike AGPD and SCAn, VBD and ABL do not rely on an auxiliary dataset. The standard deviation also shows that VBD is a stable method.
Table 5: Summary of the poisoned-instance detection $F_{1}$-score (%) averaged over all the 80 experiments.
@@
4.4 Discussion on Explainability
The proposed method extracts the salient part of the attack trigger along with its binary mask. The extracted trigger provides a quick and general visual explanation of the attack. Moreover, the binary mask can be used to isolate the pixels responsible for the malicious behavior of an instance detected as poisoned. Figure 5 illustrates the explainability of the method on a blended attack on Tiny ImageNet, using a $3\times 3$ white square trigger (shown in subfigure 5a) with a blending factor of $50\%$. The VBD method computes both the trigger and its corresponding mask, shown in subfigures 5d and 5e, respectively. The extracted pattern approximates the malicious trigger. Although the method cannot fully recover the original colors of the trigger due to the blending, the result remains sufficiently expressive to identify and explain the attack. The generated mask can then be applied to any instance detected as poisoned to isolate the malicious pixels and explain the VBD decision. The instance in subfigure 5c is detected by VBD as poisoned. The extracted mask in subfigure 5e is then used to localize the malicious pixels, as shown in subfigure 5f, where these pixels are outlined in red.
Refer to caption Refer to caption Refer to caption
(a) trigger (b) clean (c) malicious
Refer to caption Refer to caption Refer to caption
(d) extracted trigger (e) extracted mask (f) malicious pixels
Figure 5: Illustration of the method’s explainability on a blended attack using a white square trigger (a) with a blending factor of 0.5. The pixels responsible for the malicious misclassification are outlined in red in (f). In (a) and (d), transparency is represented by the gray background.